# Eddington envelopes: The fate of stars on parabolic orbits tidally disrupted by supermassive black holes

### v1

@@ Abstract
Stars falling too close to massive black holes in the centres of galaxies can be torn apart by the strong tidal forces. Simulating the subsequent feeding of the black hole with disrupted material has proved challenging because of the range of timescales involved. Here we report a set of simulations that capture the relativistic disruption of the star, followed by one year of evolution of the returning debris stream. These reveal the formation of an expanding asymmetric bubble of material extending to hundreds of astronomical units — an Eddington envelope with an optically thick inner region. Such envelopes have been hypothesised as the reprocessing layer needed to explain optical/UV emission in tidal disruption events, but never produced self-consistently in a simulation. Our model broadly matches the observed light curves with low temperatures, faint luminosities, and line widths of 10,000–20,000 km/s.
@@ 1 Introduction
The wider problem is that few calculations exist that follow the debris from disruption to fallback for a parabolic orbit with the correct mass ratio. The challenge is to evolve a main-sequence star on a parabolic orbit around a SMBH from disruption and to follow the subsequent accretion of material (Metzger & Stone, 2016). The dynamic range involved when a star on a parabolic orbit is tidally disrupted by a SMBH is greater than four orders of magnitude: the tidal disruption radius is 50 times the gravitational radius, where general relativistic effects are important, while the apoapsis of even the most bound material is another factor of 200 further away. This challenge led previous studies to consider a variety of simplifications (Stone et al., 2019): i) reducing the mass ratio between the star and the black hole by considering intermediate mass black holes (Ramirez-Ruiz & Rosswog, 2009; Guillochon et al., 2014); ii) using a Newtonian gravitational potential (Evans & Kochanek, 1989; Rosswog et al., 2008; Lodato et al., 2009; Guillochon et al., 2009; Golightly et al., 2019) or post-Newtonian approximations (Ayal et al., 2000; Hayasaki et al., 2013, 2016); iii) simulating only the first passage of the star (Evans & Kochanek, 1989; Laguna et al., 1993; Khokhlov et al., 1993; Frolov et al., 1994; Diener et al., 1997; Kobayashi et al., 2004; Guillochon et al., 2009; Guillochon & Ramirez-Ruiz, 2013; Tejeda et al., 2017; Gafton & Rosswog, 2019; Goicovic et al., 2019); and iv) assuming stars initially on bound, highly eccentric orbits instead of parabolic orbits (Sadowski et al., 2016; Hayasaki et al., 2016; Bonnerot et al., 2016; Liptai et al., 2019; Hu et al., 2024).
@@ 2 Methods
We modelled the star as a polytropic sphere with polytropic index using particles at our highest resolution and approximately K particles in our medium resolution calculations (note that real stars are expected to be more concentrated, which may alter the fallback rate, delaying the peak with respect to a polytropic model; e.g. Law-Smith et al. 2019), and include the perturbation in the metric due to Newtonian self-gravity in our simulations to hold the star together during the pre-disruption phase.
@@ 2.1 Synthetic lightcurves
where is the intensity, is the optical depth and is the frequency. For simplicity we assumed local thermodynamic equilibrium such that , where is the Planck function. We also neglected time-dependent and General Relativistic effects and special relativistic corrections to the shape of the photosphere (Abramowicz et al., 1991). That is, we considered rays to travel in a straight line towards the observer. We solved for the optical depth using
where and and is the primitive density (in the local frame). The correction factor accounts for the optical depth change caused by the moving photosphere (Abramowicz et al., 1991; Ogura & Fukue, 2013). We also computed in the observer’s frame (i.e. using ), accounting for the Doppler shift due to the motion of the photosphere.
We found the relativistic corrections to the optical depth to be negligible in practice (because our typical ), giving a 1-2% correction to the total luminosity. This also justifies our neglect of the photospheric shape distortions. A potentially more significant effect is the difference between the ‘thermalisation surface’ and the photosphere in scattering-dominated flows (Abramowicz et al., 1991; Ogura & Fukue, 2013). Full treatment of this issue is beyond the scope of our paper (requiring a full 3D solution of the radiative transfer problem including scattering), but implies that we may underestimate the true temperature at the photosphere.
@@
where is the mean molecular weight and is the radiation constant. We solved this for each particle via Newton-Raphson iteration until .
We assumed the opacity to be dominated by electron scattering, i.e. , where is the Thomson cross section and we computed by solving in a simple way for the Hydrogen ionisation fraction, from the Saha equation (assuming H only for simplicity) by solving the quadratic
where is the Hydrogen number density. In the limit of full ionisation this gives cm2/g. The main effect of (4) is to make regions with K transparent.
We assumed a grid of pixels in the image plane with one ray per pixel, integrating back to front, each pixel containing 128 frequency bins spaced evenly in between and Hz. Since our data consist of SPH particles, we performed the ray trace by sorting the particles by (the line of sight towards the observer, not necessarily the original ), considering the propagation of the ray bundle in turn through each particle, computing the optical depth integral (2) through each particle using the smoothing kernel as described in Price (2007). We thus obtained a map of on our grid of pixels at 128 different frequencies. To ‘observe’ the flow from different lines of sight, we simply rotated the particle distribution around the , or axes while keeping the location of the observer fixed at in the rotated coordinates (, and ).
Synthetic spectra were computed by integrating over all pixels in the image. To compute the bolometric luminosity we integrated the resulting spectrum over all frequencies using the trapezoidal rule on the logarithmic frequency grid. To test our procedure we confirmed that imaging a sphere of gas modelled with each particle given a uniform temperature K and a constant opacity produced a luminosity equal to the solar luminosity.
@@ 3 Results
Figure 2 shows the density (top), radial velocity (middle) and temperature distributions (bottom) of particles a year after disruption, with temperature calculated from Equation (3). The density of the Eddington envelope is g cm-3, approximately six orders of magnitude lower than the density of the stream, which is g cm-3. The typical outflow velocity is 104 km s-1. After one year there is approximately 67% ( M⊙) of the stellar mass in the incoming or unbound stream, 31.2% ( M⊙) in the envelope, 0.8% ( M⊙) in the eccentric disc, while 1% ( M⊙) has been accreted by the black hole.
The top panel of Figure 3 shows the energy dissipation rate in our simulation, . This was computed by computing the time derivative of the thermal energy caused by irreversible heating terms due to viscosity and shocks. We find that it is dominated by regions close the accretion radius . The energy dissipation rate rises rapidly from 26 days, when accretion begins, until 50 days after disruption when it reaches a peak of erg s-1 (roughly the Eddington limit). This is followed by a power law decay that does approximately match the expected energy dissipation from the mass fallback rate (Rees, 1988; Phinney, 1989), albeit scaled down by approximately two orders of magnitude. We compare this to the luminosity derived from our measured mass accretion rate , given by . Here, is the efficiency based on our inner accretion radius. However, since our is inside the last stable orbit, matter crossing this radius is mostly advected into the black hole rather than radiated, so converting it into a luminosity simply measures the amount of energy that would have been dissipated to reach this radius if all the matter were on a perfectly circular orbit. The resulting ‘luminosity’ follows the same trend as the actual energy dissipation rate above, but is roughly 0.5–1 orders of magnitude greater throughout the simulation. This suggests that there is some material accreted almost radially, with little to no orbital energy dissipated before accretion (Svirski et al., 2017).
@@
Figure 4 shows the temperature at the ‘photosphere’ of the outflowing envelope, produced by ray-tracing through our simulation as described in Appendix C. An expanding emission region of 10–100 au is consistent with those inferred from TDE observations (van Velzen et al., 2021). Photospheric temperatures are in the range expected to produce blackbody emission at UV/optical wavelength. Figure 2 shows that temperatures would remain in the 104–105 K range even accounting for scattering from deeper layers (see Appendix C).
Figure 5 compares the corresponding time evolution of our measured , and (bolometric luminosity, blackbody radius and temperature of the blackbody fit to optical wavelengths, respectively) along six different lines of sight (solid lines) to the evolution of observed TDEs inferred from black body fits taken from van Velzen et al. (2021) and Neustadt et al. (2020) (dots). The equivalent plot for the more penetrating encounter is shown on the right.
@@
The photospheric radius (centre panels), when computed from fitting to the synthetic spectra in the optical band (Appendix C), is relatively insensitive to the line-of-sight, giving typical radii of 10–100 au (– cm), in the same range as the observations. The photospheric radius reaches a peak after roughly 150–200 days in the calculation, and days in the calculation, the latter more closely matching the typical day rise time seen in most of the observations. On the other hand some TDE candidates such as ASASSN-18jd (Neustadt et al., 2020) (see legend) do show a more or less flat luminosity evolution over a 1 yr timescale, similar to our lightcurves. The transients AT2018hco and AT2019iih show the closest match to the photospheric radius evolution in our calculation, particularly for days.
@@ Discussion
Once accretion commences in the optically thick environment with inefficient cooling (we assumed adiabatic gas), the gravitational energy of accreting material launches mechanical outflows, yielding a large, asymmetric, optically thick outflowing Eddington envelope (Fig. 4).
@@
That our Eddington envelope, or ‘Black Hole Sun’, is optically thick is demonstrated by our estimated photon diffusion timescales of 10–100 days shown in the bottom panel of Figure 3. This means that the debris produced by the disruption cannot cool efficiently, leading to the formation of a photosphere. Synthetic spectral energy distributions (Appendix C) show this leads to ratios of optical to X-ray emission of order –, similar to those observed (van Velzen et al., 2021).
@@ 4 Conclusions
In summary, we simulated the tidal disruption of 1M⊙ polytropic stars around a solar mass black hole, using general relativistic hydrodynamics in the Kerr metric, following the simulations for one year post disruption. Our main finding, similar to our recent finding for real stars on eccentric orbits (Hu et al., 2024) is the production of an optically thick ‘Eddington envelope’ hypothesised by a number of previous authors (e.g. Loeb & Ulmer, 1997; Metzger & Stone, 2016).
Our calculations predict:
@@ Appendix A Radiatively efficient cooling and disc formation
Figure 6 demonstrates that disc formation can occur more rapidly if the gas is allowed to cool. We performed a medium-resolution simulation with the same parameters as in Fig. 1 but using an isentropic approximation in which we discard all irreversible heating due to viscosity and shocks, but retain heating and cooling due to work; see Liptai et al. 2019). This shows that a compact, circularised disc can more efficiently form if the stream remains cold. The formation process is similar to the stream-stream collisions seen in simulations of tidally disrupted stars on bound orbits (Hayasaki et al., 2016; Bonnerot et al., 2016) caused by the general relativistic orbital precession (Rees, 1988), except that the collisions occur at relatively large distances from the black hole (s of au) for our encounter.
The second panel of Figure 6 quantifies the degree of relativistic apsidal precession, plotting three sample trajectories computed by integrating the geodesic equations for three sample particles in the returning debris stream with (solid lines), which may be compared to the corresponding Newtonian trajectories (dashed lines). The relativistic precession, while small for , nevertheless leads to self-intersection of the stream and hence disc formation.
Figure 7 shows the same calculation but performed with a spinning black hole (; ). The disc formation process can be seen to occur in a similar manner. The second panel shows the same set of orbital trajectories as in Figure 6, but computed with . These are indistinguishable from those shown in Figure 6, indicating that the spin does not play a large role in the disc formation process for encounters. The main difference is that in the case of a spinning black hole, the disc that is subsequently formed undergoes differential Lense-Thirring precession leading to disc tearing as predicted by (Nixon et al., 2012; Nealon et al., 2015). Such differentially precessing rings are thought to be responsible for Quasi-Periodic Oscillations in black hole accretion discs (Stella & Vietri, 1998) and are also observed in tidal disruption events (Pasham et al., 2019).

### v2

@@ Abstract
Stars falling too close to massive black holes in the centres of galaxies can be torn apart by the strong tidal forces. Simulating the subsequent feeding of the black hole with disrupted material has proved challenging because of the range of timescales involved. Here we report a set of simulations that capture the relativistic disruption of the star, followed by one year of evolution of the returning debris stream. These reveal the formation of an expanding asymmetric bubble of material extending to hundreds of astronomical units — an outflowing Eddington envelope with an optically thick inner region. Such outflows have been hypothesised as the reprocessing layer needed to explain optical/UV emission in tidal disruption events, but never produced self-consistently in a simulation. Our model broadly matches the observed light curves with low temperatures, faint luminosities, and line widths of 10,000–20,000 km/s.
@@ 1 Introduction
The wider problem is that few calculations exist that follow the debris from disruption to fallback for a parabolic orbit with the correct mass ratio. The challenge is to evolve a main-sequence star on a parabolic orbit around a SMBH from disruption and to follow the subsequent accretion of material (Metzger & Stone, 2016). The dynamic range involved when a star on a parabolic orbit is tidally disrupted by a SMBH is greater than four orders of magnitude: the tidal disruption radius is 50 times the gravitational radius, where general relativistic effects are important, while the apoapsis of even the most bound material is another factor of 200 further away. This challenge led previous studies to consider a variety of simplifications (Stone et al., 2019): i) reducing the mass ratio between the star and the black hole by considering intermediate mass black holes (Ramirez-Ruiz & Rosswog, 2009; Guillochon et al., 2014); ii) using a Newtonian gravitational potential (Evans & Kochanek, 1989; Rosswog et al., 2008; Lodato et al., 2009; Guillochon et al., 2009; Golightly et al., 2019), pseudo-Newtonian (Hayasaki et al., 2013; Bonnerot et al., 2016) or post-Newtonian approximations (Ayal et al., 2000; Hayasaki et al., 2016); iii) simulating only the first passage of the star (Evans & Kochanek, 1989; Laguna et al., 1993; Khokhlov et al., 1993; Frolov et al., 1994; Diener et al., 1997; Kobayashi et al., 2004; Guillochon et al., 2009; Guillochon & Ramirez-Ruiz, 2013; Tejeda et al., 2017; Gafton & Rosswog, 2019; Goicovic et al., 2019); and iv) assuming stars initially on bound, highly eccentric orbits instead of parabolic orbits (Sadowski et al., 2016; Hayasaki et al., 2013, 2016; Bonnerot et al., 2016; Liptai et al., 2019; Hu et al., 2024).
@@ 2 Methods
We modelled the star as a polytropic sphere with polytropic index using equal mass particles at our highest resolution and approximately K particles in our medium resolution calculations (real stars are expected to be more concentrated, which may alter the fallback rate, delaying the peak with respect to a polytropic model; e.g. Law-Smith et al. 2019), and include the perturbation in the metric due to Newtonian self-gravity in our simulations to hold the star together during the pre-disruption phase.
For our main simulations we employed an adiabatic approximation, implying that energy is trapped or advected rather than radiated. Appendix A reports a set of isentropic calculations where we assume shock heating to be immediately radiated away. The adiabatic simulations capture all accretion feedback down to inside which particles are assumed to accrete without radiating.
@@ 2.1 Synthetic lightcurves
where is the intensity, is the optical depth and is the frequency. For simplicity we assumed local thermodynamic equilibrium such that , where is the Planck function. We also neglected time-dependent and General Relativistic effects and special relativistic corrections to the shape of the photosphere (Abramowicz et al., 1991). That is, we considered rays to travel in a straight line to the observer. We solved for the optical depth using
where and and is the density in the rest frame of the fluid. The correction factor accounts for the optical depth change caused by the moving photosphere (Abramowicz et al., 1991; Ogura & Fukue, 2013). We also computed in the observer’s frame (i.e. using ), accounting for the Doppler shift due to the motion of the photosphere.
We found the relativistic corrections to the optical depth to be negligible in practice (because our typical ), giving a 1-2% correction to the total luminosity. This also justifies our neglect of photospheric shape distortions. Potentially more significant is the difference between the ‘thermalisation surface’ and the photosphere in scattering-dominated flows (Abramowicz et al., 1991; Ogura & Fukue, 2013). Full treatment of this issue is beyond the scope of our paper (requiring a full 3D solution of the radiative transfer problem including scattering), but implies that we may underestimate the true temperature at the photosphere.
@@
where is the mean molecular weight and is the radiation constant. We solved this for each particle via Newton-Raphson iteration until , where the initial guess ].
We adopted the opacity prescription from Matsumoto & Metzger (2022, see their figure 3 for a comparison to OPAL tables), where
which includes molecular opacity cm2/g where is the metallicity, opacity cm2/g, bound-free and free-free opacity via Kramer’s law cm2/g with (both formulae assuming in g/cm3 and in K) and electron scattering , where is the Thomson cross section. We computed by solving for the hydrogen ionisation fraction, from the Saha equation (assuming H only for simplicity) by solving the quadratic
where is the hydrogen number density. In the limit of full ionisation this gives cm2/g. The main effect of (5) is to make regions with K transparent. In practice we found identical results with since electron scattering dominates the opacity at our typical photospheric densities (g/cm3) and temperatures (K).
We assumed a grid of pixels in the image plane with one ray per pixel, integrating back to front, each pixel containing 128 frequency bins spaced evenly in between and Hz. A caveat is that (4) assumes grey opacities but we add blackbody spectra in a frequency-dependent manner. This is irrelevant anyway since the dominant opacity is grey. Since our data consist of SPH particles, we performed the ray trace by sorting the particles by (the line of sight towards the observer, not necessarily the original ), considering the propagation of the ray bundle in turn through each particle, computing the optical depth integral (2) through each particle using the smoothing kernel as described in Price (2007). We thus obtained a map of on our grid of pixels at 128 different frequencies. To ‘observe’ the flow from different lines of sight, we simply rotated the particle distribution around the , or axes while keeping the location of the observer fixed at in the rotated coordinates (, and ). At early times the photosphere can be unresolved by the SPH particles (e.g. Hatfull et al., 2021). To check for this we flagged pixels in our synthetic images where a single particle contributed more than while . We then set a criterion on the unresolved area compared to the total area where . When of pixels in the photosphere are poorly resolved we show the lightcurve using semi-transparent lines.
Synthetic spectral energy distributions were computed by integrating over all pixels in the image. Inferred bolometric luminosities were then computed by integrating the blackbody spectrum fitted to the optical band. To test our procedure we confirmed that imaging a sphere of gas modelled with each particle given a uniform temperature K and a constant opacity produced a luminosity equal to the solar luminosity.
@@ 3 Results
Figure 2 shows the density (top), temperature (middle) and radial velocity (bottom) distributions of particles a year after disruption, with temperature calculated from Equation (3). The density of the outflowing envelope is g cm-3, approximately six orders of magnitude lower than the density of the stream, which is g cm-3. The typical outflow velocity is 104 km s-1. After one year there is approximately 67% ( M⊙) of the stellar mass in the incoming or unbound stream, 31.2% ( M⊙) in the envelope, 0.8% ( M⊙) in the eccentric disc, while 1% ( M⊙) has been accreted by the black hole.
The top panel of Figure 3 shows the energy dissipation rate in our simulation, . This was computed by computing the time derivative of the thermal energy caused by irreversible heating terms due to viscosity and shocks. We find that it is dominated by regions close the accretion radius . The energy dissipation rate rises rapidly from 26 days, when accretion begins, until 50 days after disruption when it reaches a peak of erg s-1 (roughly the Eddington limit). This is followed by a power law decay that does approximately match the expected energy dissipation from the mass fallback rate (Rees, 1988; Phinney, 1989), but with lower than the mass fallback rate by approximately two orders of magnitude. We compare this to the luminosity derived from our measured mass accretion rate , given by . Here, is the efficiency based on our inner accretion radius. However, since our is inside the last stable orbit, matter crossing this radius is mostly advected into the black hole rather than radiated, so converting it into a luminosity simply measures the amount of energy that would have been dissipated to reach this radius if all the matter were on a perfectly circular orbit. The resulting ‘luminosity’ follows the same trend as the actual energy dissipation rate above, but is 0.5–1 orders of magnitude greater throughout the simulation. This suggests some material is accreted almost radially, with little to no orbital energy dissipated before accretion (Svirski et al., 2017).
@@
Figure 4 shows the temperature at the ‘photosphere’ of the outflowing envelope, produced by ray-tracing through our simulation as described in Sec. 2.1. An expanding emission region of 10–100 au is consistent with those inferred from TDE observations (van Velzen et al., 2021). Photospheric temperatures are in the range expected to produce blackbody emission at UV/optical wavelength. Figure 2 shows that temperatures would remain in the 104–105 K range even accounting for scattering from deeper layers (see Appendix C).
Figure 5 compares the corresponding time evolution of our measured , and (inferred bolometric luminosity, blackbody radius and temperature of the blackbody fit to optical wavelengths, respectively) from six different lines of sight (solid lines) to the evolution of observed TDEs inferred from black body fits taken from van Velzen et al. (2021) and Neustadt et al. (2020) (dots). The equivalent plot for the more penetrating encounter is shown on the right.
@@
The photospheric radius (centre panels), when computed by fitting to synthetic spectra in the optical band (Appendix C), is relatively insensitive to the line-of-sight, giving typical radii of 10–100 au (– cm), in the same range as the observations. The photospheric radius reaches a peak after roughly 150–200 days in the calculation, and days in the calculation, the latter more closely matching the typical day rise time seen in most of the observations. On the other hand some TDE candidates such as ASASSN-18jd (Neustadt et al., 2020) (see legend) do show a more or less flat luminosity evolution over a 1 yr timescale, similar to our lightcurves. The transients AT2018hco and AT2019iih show the closest match to the photospheric radius evolution in our calculation, particularly for days.
@@ Discussion
Once accretion commences in the optically thick environment with inefficient cooling (we assumed adiabatic gas), the gravitational energy of accreting material launches mechanical outflows, yielding a large, asymmetric, optically thick outflowing envelope (Fig. 4).
@@
That our outflowing envelope, or ‘Black Hole Sun’, is optically thick is demonstrated by our estimated photon diffusion timescales of 10–100 days shown in the bottom panel of Figure 3. This means that the debris produced by the disruption cannot cool efficiently, leading to the formation of a photosphere. Synthetic spectral energy distributions (Appendix C) show this leads to ratios of optical to X-ray emission of order –, similar to those observed (van Velzen et al., 2021).
@@ 4 Conclusions
In summary, we simulated the tidal disruption of 1M⊙ polytropic stars around a solar mass black hole, using general relativistic hydrodynamics in the Kerr metric, following the simulations for one year post disruption. Our main finding, similar to our recent finding for real stars on eccentric orbits (Hu et al., 2024) is the production of an optically thick ‘Eddington envelope’ hypothesised by previous authors (e.g. Loeb & Ulmer, 1997; Metzger & Stone, 2016). Our calculations predict:
@@ Appendix A Radiatively efficient cooling and disc formation
Figure 6 demonstrates that disc formation can occur more rapidly if the gas is allowed to cool. We performed a medium-resolution simulation with the same parameters as in Fig. 1 but using an isentropic approximation in which we discard all irreversible heating due to viscosity and shocks, but retain heating and cooling due to work; see Liptai et al. 2019). This shows that a compact, circularised disc can more efficiently form if the stream remains cold. The formation process is similar to the stream-stream collisions seen in simulations of tidally disrupted stars on bound orbits (Hayasaki et al., 2016; Bonnerot et al., 2016) caused by the general relativistic orbital precession (Rees, 1988), except that the collisions occur at relatively large distances from the black hole (s of au) for our encounter. No optically thick envelope is formed in the isentropic case as the main energy source is switched off.
The second panel of Figure 6 quantifies the degree of relativistic apsidal precession, plotting three sample trajectories computed by integrating the geodesic equations111https://github.com/phantomSPH/phantom-geodesic for three sample particles in the returning debris stream with (solid lines), which may be compared to the corresponding Newtonian trajectories (dashed lines). The relativistic precession, while small for , nevertheless leads to self-intersection of the stream and hence disc formation.
Figure 7 shows the same calculation but performed with a spinning black hole (; ). The disc formation process can be seen to occur in a similar manner. The second panel shows the same set of orbital trajectories as in Figure 6, but computed with . These are indistinguishable from those shown in Figure 6, indicating that the spin does not play a large role in the disc formation process for encounters. The main difference is that in the case of a spinning black hole, the disc that is subsequently formed undergoes differential Lense-Thirring precession leading to disc tearing as predicted by Nixon et al. (2012) and Nealon et al. (2015). Such differentially precessing rings are thought to be responsible for Quasi-Periodic Oscillations in black hole accretion discs (Stella & Vietri, 1998) and are also observed in tidal disruption events (Pasham et al., 2019).
@@ Appendix E Effect of resolution
Conservation properties are well preserved by the numerical scheme at all resolutions. For example, in the calculation shown in Figure 1 the energy error is up to days when accretion commences, after which the total energy rises steadily due to the removal of accreted mass with negative specific energy from the simulation. Total angular momentum is conserved to (0.02%) over the entire 365 days of the simulation.